how many gallons of a 70% antifreeze solution must be mixed with 100 gallons of 25% antifreeze to get a mixture that is 60% antifreeze? Use the six-step method

Answer by checkley71(8403)

.7X+.25*100=.6(100+X)
.7X+25=60+.6X
.7X-.6X=60-25
.1X=35
X=35/.1
X=350 GALLONS OF 70% ANTIFREEZE WILL BE NEEDED.
PROOF
.7*350+25=.6*450
245+25=270
270=270